N is 80% of P
N to the root of 2 is what percent of P to the root of 2?

Answer by Theo(671)

n = 80% of p means that:

n = .8*p

take the square root of both sides of this equation and you get:

sqrt(n) = sqrt(.8*p) = sqrt(.8) * sqrt(p)

sqrt(.8) = .894427191 * 100% = 89.4427191%

your answer is:

sqrt(n) = 89.4427191% of sqrt(p)

to confirm, let n = 100

this makes p = 100/.8 = 125

equation of

n = .8*p becomes

100 = .8*125

take sqrt of both sides to get:

sqrt(100) = sqrt(.8*125) = sqrt(.8) * sqrt(125)

this becomes:

10 = .894427191 * 11.18033989

this becomes:

10 = 10

Answer is confirmed.